OpenStudy (anonymous):

Find the quotient and simplify. \frac{ (y-9)^2 }{ 2}div \frac{ 2y-18 }{ 4 }

OpenStudy (anonymous):

\[\frac{ (y-9)^2 }{2}\div \frac{ 2y-18 }{ 4 }\]

OpenStudy (kropot72):

The expression can be rewritten as \[\frac{(y-9)^{2}}{2}\times \frac{4}{2(y-9)}\] Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Good! So what is the result if you multiply the two numerators and the two denominators of the rewritten expression, using this rule: \[\frac{a}{b}\times \frac{c}{d}=\frac{ac}{bd}\]

OpenStudy (anonymous):

y^3+1458/8

OpenStudy (kropot72):

No. You tried multiplying the numerators and the denominators of the original question. I asked you to multiply the two numerators and the two denominators of the REWRITTEN expression. I gave you the rewritten expression in my first posting.

OpenStudy (kropot72):

@corybory Do you understand?